OpenStudy (anonymous):

(x-6)/)(x+5) inverse function?

OpenStudy (anonymous):

\[y=\frac{x-6}{x+5}\] like that?

OpenStudy (anonymous):

switch x and y because that is what the inverse does, get \[x=\frac{y-6}{y+5}\] and solve for \(y\) with a bunch of algebra

OpenStudy (anonymous):

given f(x)=(x-6)/(x+5), i have to find f^-1(x)=

OpenStudy (anonymous):

start by clearing the fraction to get \[x(y+5)=y-6\]\[xy+5x=y-6\] then put everything with a \(y\) in it on one side, everything else on the other and get \[y-xy=5x+6\]

OpenStudy (anonymous):

factor on the left to get \[(1-x)y=5x+6\] then divide and get \[y=\frac{5x+6}{1-x}\]

OpenStudy (anonymous):

that one is your inverse

OpenStudy (anonymous):

I see now!! thanks.

OpenStudy (anonymous):

yw